A STUDENT DEPOSIT 1500.00 IN A 9% ACCOUNT TODAY HE INTENDS TO DEPOSIT ANOTHER 3000.00 AT THE END OF TWO YEARS. HE PLAN TO PURCHASE IN 5 YEARS HIS FAVORITE SHOES WORTH OF 5000.00. CALCULATE THE MONEY THAT WILL BE LEFT IN HIS ACCOUNT ONE YEAR AFTER THE PURCHASE
Found 2 solutions by mananth, rfer:
Answer by mananth(12270)

$1500 ---- 5 years--------9%
=1500*5*0.09= 675
$3000-3years - 9%
3000*3*0.09=810
Total interest earned = 810+675= 1485
Total in bank after 5 years = 5000+1485= 6485
Less cost of shoes $5000
=1485.
he will earn 9% on 1485 for 1 year
Interest= 1485*0.09*1
=133.65
Money in Bank after 6 years = 1080 + 133.65=1177.20
=$1213.65

Answer by rfer(12676)

A=1500(1.09)^2
A=1782.15
-----------
A=(1782.15+3000)(1.09)^3
A=4782.15(1.295)
A=6193
--------------------
a=6193-5000=1193
A=1193(1.09)
A=$1300.37